OpenStudy (anonymous):

which of the following tables shows the correct steps to transform x^2+8x+15=0 into the form of (x-p)^2=q (p and q are integers)

OpenStudy (anonymous):

table 1: step 1.x^2+8x+15-1=0-1 step 2.x^2 +8x+14=-1 step 3.(x+4)^2=-1

OpenStudy (anonymous):

table 2: step 1.x^2+8x+15-2=0-2 step 2.x^2+8x+13=-2 step 3.(x+4)^2-2

OpenStudy (anonymous):

table 3: step 1.x^2+8x+15+1=0+1 step 2.x^2+8x+16=1 step 3.(x+4)^2=1

OpenStudy (anonymous):

table 4: step 1.x^2+8x+15+1=0+2 step 2.x^2+8x+17=2 step 3.(x+4)^2=2

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@poopsiedoodle

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

lets before see can u convert x^2+8x+15=0 into the form of (x-p)^2=q ?

OpenStudy (ikram002p):

x^2+8x+15=0 is quadratic equation we can convert it through اComplete square

OpenStudy (ikram002p):

so the genera form for quadratic equation is ax^2+bx+c=0 first step to convert is add (b/2)^2 for both side second step the equation would be ax^2+bx+(b/2)^2=(b/2)^2-c third step convert to (x+(or -) p/2)^2=(b/2)^2-c (x+q)^2=p :D

OpenStudy (ikram002p):

ur equation is x^2+8x+15=0 x^2+8x=-15 so a=1 b=8 c=15 first step add (b/2)^2 for both side"(b/2)^2=(8/2)^2=4^2) x^2+8x + 4^2=-15 + 16 second step do the arithmetic x^2+8x+16=1 step 3 convert (x+4)^2=1 p=-4 q=1

OpenStudy (ikram002p):

hope u got it :D

OpenStudy (anonymous):

so table 3?

OpenStudy (anonymous):

@ikram002p

OpenStudy (ikram002p):

yess